Grade 12Analytical Geometry

Tips For Solving Coordinate Geometry From Arihant Sk Goyal

Saif Coolx
15 Years agoGrade 12

3 Answers

SAGAR SINGH - IIT DELHI
15 Years ago

Dear saif,

First of all learn the basic concepts and after your concepts are clear then you can solve the solved examples and once you have enuf confidence you can go for Arihant book...

 

 

 

Please feel free to post as many doubts on our discussion forum as you can.

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 

win exciting gifts by answering the questions on Discussion Forum



rex gautam
15 Years ago

1.what does equation (a-b)(x^2+y^2)-2abx=0 become if the origin be moved to the points [ab/a-b,0]?

2.Ther equation x^2+2xy+4=0 is transformed to the parallel axes through the point (6,m).For what value of m its new origin form passes through the new origin?

3.Transform the equation x^2-3xy+11x-12y+36=0 to parallel then the axes through the point (-4,1) becomes ax^2+bxy+1=0 then b^2-a=?

Anu
7 Years ago
Ther equation x^2+2xy+4=0 is transformed to the parallel axes through the point (6,m).For what value of m its new origin form passes through the new origin?